OpenStudy (lollylau):

[Easy] Peter and John need to use 6 days to complete some work. John and Lucas need 9 days to finish it, and Peter and Lucas need 15 days to finish it. How many days do they have to use if they do the work together?

OpenStudy (lollylau):

anyone?

OpenStudy (lollylau):

They're still busy doing the previous one. "sigh"

OpenStudy (anonymous):

this is a set of diophantine equations right

OpenStudy (anonymous):

You can do it using time and work formula.

OpenStudy (lollylau):

get three people named Peter, John and Lucas and start :)

OpenStudy (anonymous):

work = work/time * time?

OpenStudy (anonymous):

alright in Python I will create Peter, John, and Lucas objects and see how long they take to do the thing

OpenStudy (anonymous):

Do you need to use python for EVERYTHING?

OpenStudy (lollylau):

easy solution below.

OpenStudy (lollylau):

Peter = p John = j Lucas = l p+j = 1/6 j+l = 1/9 p+l = 1/15.

OpenStudy (lollylau):

add p, j, and l and find the reciprocal.

OpenStudy (anonymous):

Yeah, I knew that, just too lazy to work it out.

OpenStudy (anonymous):

yea it is a bunch of arithmetic \[\frac{180}{31}\] i think

OpenStudy (lollylau):

nice! you got it correct.

OpenStudy (mertsj):

Of course he did.

OpenStudy (anonymous):

my arithmetic somethings fails me though

OpenStudy (lollylau):

namely careless.